# Keeping in Place After the Storm—Emergency Assistance and Evictions

### v2

@@ 4 Conclusion
Using granular data on evictions and loan-level data on payday loans, we provide evidence of disaster induced increase in evictions and reliance on high cost credit in the state of Florida following major storms and hurricanes. We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit. Overall, this suggests that major household level impacts such as access to housing are sensitive to the availability of emergency liquidity, suggesting the utility for such programs and the need to better understand their availability in a world subject to an increasing propensity of climate shock events.

### v3

@@ 4 Conclusion
Using granular data on evictions and loan-level data on payday loans, we provide evidence of disaster induced increase in evictions and reliance on high cost credit in the state of Florida following major storms and hurricanes. We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit. Overall, this suggests that major household level impacts such as access to housing are sensitive to the availability of emergency liquidity, suggesting the utility for such programs while highlighting the need to better understand administrative cutoffs to access.